# Model Predictive and Reinforcement Learning Methods for Active Flow Control of an Airfoil with Dual-point Excitation of Plasma Actuators

### v1

@@ 3.4 Numerical method
The computational fluid dynamics (CFD) simulations in this study were conducted using a pressure-based finite volume solver. The simulations were transient and performed in a two-dimensional planar space, with gravity effects neglected.
The air with constant properties, a density of 1.225 kg/m3 and a viscosity of $1.7894\times 10^{-5}$ kg/($\cdots$), was used in the simulation. A coupled scheme with Rhie-Chow momentum-based flux type [37] was implemented for pressure and velocity fields. To achieve high accuracy in the numerical solution, least squares cell-based gradient discretization was applied. Second-order discretization was used for pressure, while bounded central differencing was applied for momentum. Second-order upwind schemes were used for both turbulent kinetic energy and specific dissipation rate.
The transient nature of the flow was captured using a bounded second-order implicit scheme, which provides a balance between stability and accuracy for time-dependent simulations. This setup was carefully selected to provide a comprehensive and accurate representation of the flow dynamics over the airfoil with active flow control using plasma actuators.
@@ 4.1 Adaptive MPC
where polynomials $E_{j}$ and $F_{j}$ are uniquely defined for a given $\tilde{A}$ that have degrees of $j-1$ and $n_{a}$, respectively. If eq. (8) is multiplied by $\Delta E_{j}(z^{-1})z^{j}$ and using eq. (11), we can find the $j$-step ahead prediction of the system output:
@@
where $G_{j}(z^{-1})=E_{j}(z^{-1})B(z^{-1})$. Thus, the degree of polynomial $E_{j}(z^{-1})$ equals $j-1$, the noise term $e(t+j)$ in equation eq. (12) is all in the future. Therefore the best prediction of $y(t+j)$ is:
@@
The parameters required for implementing the Adaptive GPC algorithm are described in TABLE. 1. These parameters are used in the results.
Table 1: Parameter values for adaptive GPC algorithm.
@@ 4.2 TDRL
where, $s$ is the current state from which the action $a$ is taken, $s^{\prime}$ is the new state resulting from taking action $a$, $\alpha$ is the learning rate, $r$ is the reward received after taking action $a$ in state $s$, $\gamma$ is the discount factor, $Q(s,a)$ is the current Q-value for taking action $a$ in state $s$, and $\max_{a^{\prime}}Q(s^{\prime},a^{\prime})$ is the maximum Q-value among all possible actions from the new state $s^{\prime}$. The algorithm decides whether to explore new actions or exploit the current knowledge by comparing a random number with the exploration rate ($\epsilon$). $\epsilon$ is decayed by multiplying it by exploration decay ($\kappa$) in every controller time step. The parameters required for implementing the TDRL algorithm are described in TABLE. 2.
Table 2: Parameter values for TD and DQL algorithms.
@@
TimeStep $1\times 10^{-4}$ $1\times 10^{-4}$ Simulation time step
@@ 4.3 Deep Q - Learning
DQL is a model-free RL algorithm that combines Q-Learning with deep neural networks. It is designed to handle environments with high-dimensional state spaces where traditional Q-Learning would be infeasible [45]. DQL approximates the Q-value function using a neural network, which takes the current state as input and outputs Q-values for each possible action. The algorithm iteratively updates the Q-values based on the observed rewards and transitions. Like the previous method, rewards and states are calculated based on the environment’s responses to the agent’s actions. The Q-value update rule is given by:
$Q(s,a)\leftarrow Q(s,a)+\alpha\left[r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime})-Q(s,a)\right]$ (25)
Where $\alpha$ is the learning rate that determines the extent to which the new information overrides the old information. If $\alpha=1$, the equation simplifies to:
@@
Figure 5: DQL Network Architecture. W and B indicate weight and Bias of layers, respectively.
@@ 4.4 Signal processing integrated with Deep Q-Learning
Figure 6: Signal processing integrated with DQL Network Architecture.
@@ 5.1 Baseline Flow Simulation and Validation
This subsection focuses on the baseline flow simulation and its validation against experimental data. Figure  7 demonstrates the pressure coefficient ($C_{p}$) distribution along the airfoil surface against experimental data of Mallor et al. [47] at an angle of attack $\alpha=14^{\circ}$, for a Reynolds number $Re=4.0\times 10^{5}$, which shows a satisfactory agreement. Furthermore, the dimensionless wake velocity profile at T.E. is compared with other experimental and numerical studies of Tabatabaei et al. [48] in Fig. 8. In this figure, $U_{e}$ is the edge velocity of the wake. The results indicate that the present simulation follows the overall trend of the numerical and experimental data.
Since this study aims to investigate the stall condition, the rest of the paper is focused on $\alpha=15^{\circ}$. Fig. 9 presents the velocity magnitude, turbulent kinetic energy, and vorticity magnitude contours in the wake region. According to Figs.  7 and  9(a), the separation point is about 0.55c. Estimating the flow separation point location is necessary to determine the actuator placement on the suction side. According to previous research of Ebrahimi and Hajipour [32], the actuator placement upstream of the separation point is more effective in mitigating the flow separation. Figs. 9(b) and 9(c) demonstrate two shear layers following the separation point and T.E., which is the basis of utilizing the dual-point excitation technique. Thus, the pressure side actuator is placed on the airfoil’s trailing edge.
Figure 7: Comparison of $C_{p}$ distribution along the $x/c$ between experimental and CFD data.
Figure 8: Comparison of wake velocity profiles at T.E.
Figure 9: Contours of baseline flow: (a) Velocity magnitude, (b) Turbulent kinetic energy, and (c) Vorticity magnitude.
@@ 5.2 Adaptive MPC
The adaptive MPC was evaluated across two target values to assess its effectiveness in controlling flow separation. The analysis highlights the control strategy’s adaptability and impact on system stability and performance. First, a target $C_{l}$ of 1.6 was selected as the desired output. The objective was to adjust the excitation frequency of the plasma actuators in real time to achieve and maintain this lift coefficient. The control system dynamically responds to the airfoil’s transient aerodynamic characteristics to ensure that the $C_{l}$ reaches and stabilizes around the desired value of 1.6.
Figure 10: Desired $C_{l}$ of 1.6 using adaptive MPC and its corresponding excitation frequencies.
The provided graph, as shown in Fig. 10, illustrates the performance of the adaptive MPC system in achieving the desired lift coefficient. The top panel of the graph shows the time series of the $C_{l}$ values along with its moving average. The bottom panel depicts the corresponding excitation frequencies applied during the control process. The lift coefficient was stabilized at around 2.8 s after activation of the controller.
To push the boundaries of the control system, aiming to achieve higher lift coefficients, a second target $C_{l}$ of 1.62 was selected. However, reaching such a high $C_{l}$ presents significant challenges for the adaptive MPC to stabilize $C_{l}$ around 1.62.
Figure 11: Desired $C_{l}$ of 1.62 using Adaptive MPC and its corresponding excitation frequencies.
As depicted in Fig. 11, the controller could not stabilize the $C_{l}$ at the desired value of 1.62. The fluctuations in the $C_{l}$ time series indicate instability as the control system struggles to maintain the target value. This instability seems to arise because $C_{l}=1.62$ is near the physical limit for the dual plasma actuator configuration, making it difficult for the controller to exert precise control. Consequently, the control system oscillates around the target, failing to maintain the given target.
@@ 5.3 TDRL
Figure 12: Exploring the optimal excitation frequency that maximizes $C_{l}$ using TDRL.
Based on Fig. 12, the control system explored various frequencies before settling on 200 Hz as the optimal frequency for maximizing $C_{l}$. The top panel shows how $C_{l}$ evolves over time, while the bottom panel depicts the selected excitation frequencies throughout the process. By the end of the simulation, the system determined that 200 Hz was the most effective frequency for achieving the maximum $C_{l}$. Notably, after four decision-makings, the controller found the optimal frequency of 200 Hz, resulting in a mean $C_{l}$ of approximately 1.619.
@@ 5.4 DQL
Figure 13: Selection of the optimal excitation frequency for maximizing $C_{l}$ using DQL.
As shown in Fig. 13, the DQL algorithm explored various frequencies before selecting 100 Hz as the optimal excitation frequency. Despite the fluctuations during the learning phase, the system eventually converged to 100 Hz as the most effective frequency for achieving a high and stable $C_{l}$. After 11 desicion makings, the controller found the optimal frequency of 100 Hz, resulting in a mean $C_{l}$ of approximately 1.619.
@@ 5.5 Signal processing integrated with DQL
In this approach, DQL was enhanced by integrating signal processing techniques to improve the accuracy and robustness of the excitation frequency selection process. The objective was to refine the control system’s ability to identify the optimal frequency for maximizing the lift coefficient. The system eventually selected 200 Hz as the optimal frequency. After 21 desicion makings, the controller found the optimal frequency of 200 Hz, resulting in a mean $C_{l}$ of approximately 1.619.
Figure 14: Selection of the optimal excitation frequency for maximizing $C_{l}$ using signal processing integrated with DQL.
The average $C_{l}$ for 100 Hz and 200 Hz are nearly identical, which made the decision-making process challenging for the network. The reason they are similar can be seen in the Fig. 15. For 100 Hz, there is one period of “on” and “off” states, while for 200 Hz, there are two periods. However, the first “on” phase for 200 Hz has a lower maximum than the “on” phase for 100 Hz, and the second “on” phase for 200 Hz reaches a higher maximum than 100 Hz. As a result, the average of the “on” states for 200 Hz becomes nearly identical to that of 100 Hz. This behavior suggests that the DQL produces similar average $C_{l}$ compared to the signal processing integrated with DQL and TDRL cases, making it harder for the network to differentiate between the two based purely on $C_{l}$ values.
Figure 15: Comparison of results from different reinforcement learning methods for determining $C_{l}$
@@ 5.6 Comparative Analysis
Figure 16: Stabilization time of different control methods.
@@
To compare the controlled flow field with the baseline flow (as shown in Fig.9), Fig.17 presents the contours of velocity magnitude, turbulent kinetic energy, and vorticity magnitude. These results are specifically provided for the signal processing integrated with DQL method. Since the flow fields for other control methods are similar, their contours are not presented. This figure demonstrates that the proposed algorithms have successfully modified the flow field, including shear layers, wake, and separation points, compared to the baseline case. The velocity magnitude contours in Figures 9(a) and 17(a) reveal a reduction in the extent of the separation zone under the influence of the flow control method. This is further supported by the notable reduction in turbulent kinetic energy levels observed in the controlled case relative to the baseline flow (Figures 9(b) and 17(b)). Additionally, the vorticity magnitude contours in Figures 9(c) and 17(c) indicate that the separated shear layer on the airfoil’s suction side is inclined closer to the airfoil surface compared to the baseline case.
@@
Figure 17: Contours of controlled flow using signal processing integrated with DQL method: (a) Velocity magnitude, (b) Turbulent kinetic energy, and (c) Vorticity magnitude.

### v2

@@ 3.4 Numerical method
[…]
A coupled scheme with Rhie-Chow momentum-based flux type [55] was implemented for pressure and velocity fields. To achieve high accuracy in the numerical solution, least squares cell-based gradient discretization was applied. Second-order discretization was used for pressure, while bounded central differencing was applied for momentum. Second-order upwind schemes were used for both turbulent kinetic energy and specific dissipation rate. The transient nature of the flow was captured using a bounded second-order implicit scheme, which provides a balance between stability and accuracy for time-dependent simulations. The CFD solver time step was set to $10^{-4}$ s, which is sufficiently smaller than the excitation periods associated with the tested actuation frequencies (50–400 Hz). This choice ensures adequate temporal resolution to accurately capture the unsteady effects of plasma forcing. The numerical setup described in this section was carefully selected to ensure accurate and robust simulation of the flow dynamics under active flow control.
@@ 4.1 Adaptive MPC
where polynomials $E_{j}$ and $F_{j}$ are uniquely defined for a given $\tilde{A}$ that have degrees of j1j-1 and $n_{a}$, respectively. If Eq. (8) is multiplied by $\Delta E_{j}(z^{-1})z^{j}$ and using Eq. (11), we can find the $j$-step ahead prediction of the system output:
@@
where $G_{j}(z^{-1})=E_{j}(z^{-1})B(z^{-1})$. Thus, the degree of polynomial $E_{j}(z^{-1})$ equals j1j-1, the noise term $e(t+j)$ in equation Eq. (12) is all in the future. Therefore the best prediction of $y(t+j)$ is:
@@
The parameters required for implementing the Adaptive GPC algorithm are described in Table 1. These parameters are used in the results.
Table 1: Parameter values for adaptive GPC algorithm. Source: Authors’ own work.
@@ 4.2 TDRL
where, $s$ is the current state from which the action $a$ is taken, $s^{\prime}$ is the new state resulting from taking action $a$, $\alpha$ is the learning rate, $r$ is the reward received after taking action $a$ in state $s$, $\gamma$ is the discount factor, $Q(s,a)$ is the current Q-value for taking action $a$ in state $s$, and $\max_{a^{\prime}}Q(s^{\prime},a^{\prime})$ is the maximum Q-value among all possible actions from the new state $s^{\prime}$. The algorithm decides whether to explore new actions or exploit the current knowledge by comparing a random number with the exploration rate ($\epsilon$). $\epsilon$ is decayed by multiplying it by exploration decay ($\kappa$) in every controller time step. The parameters required for implementing the TDRL algorithm are described in Table 2.
Table 2: Parameter values for TD and DQL algorithms. Source: Authors’ own work.
@@
TimeStep $\times 10^{-4} 1×1$$\times 10^{-4}$ Simulation time step
@@ 4.3 Deep Q - Learning
DQL is a model-free RL algorithm that combines Q-Learning with deep neural networks. It is designed to handle environments with high-dimensional state spaces where traditional Q-Learning would be infeasible [63]. DQL approximates the Q-value function using a neural network, which takes the current state as input and outputs Q-values for each possible action. The algorithm iteratively updates the Q-values based on the observed rewards and transitions. Like the previous method, rewards and states are calculated based on the environment’s responses to the agent’s actions. The Q-value update rule is given by Eq. (24). Where $\alpha$ is the learning rate that determines the extent to which the new information overrides the old information. If $\alpha=1$, the equation simplifies to:
@@
Figure 5: DQL Network Architecture. W and B indicate weight and Bias of layers, respectively. Source: Authors’ own work.
@@ 4.4 Signal processing integrated with Deep Q-Learning
Figure 6: Signal processing integrated with DQL Network Architecture. Source: Authors’ own work.
@@ 5.1 Baseline Flow Simulation and Validation
This subsection focuses on the baseline flow simulation and its validation against experimental data. Fig. 7 presents the pressure coefficient ($C_{p}$) distribution along the airfoil surface, compared with the experimental data of Mallor et al. [65] at $14^{\circ}$ angle of attack and Reynolds number $Re=4.0\times 10^{5}$, showing satisfactory agreement. Furthermore, the dimensionless wake velocity profile at the trailing edge is compared with the experimental and numerical data of Tabatabaei et al. [66] in Fig. 8, where $U_{e}$ denotes the edge velocity of the wake. The results indicate that the present simulation captures the overall trend observed in both experimental and numerical data. As this study aims to investigate near-stall conditions, the remainder of the paper focuses on $\alpha=15^{\circ}$.
Figure 7: Comparison of $C_{p}$ distribution along the $x/c$ between experimental and CFD data. Source: Authors’ own work.
Figure 8: Comparison of wake velocity profiles at the trailing edge. Source: Authors’ own work.
Fig. 9 presents contours of different flow quantities in the baseline airfoil wake region, including velocity magnitude, turbulent kinetic energy (TKE), vorticity magnitude, and the Q-criterion. Based on the pressure coefficient distribution (Fig. 7) and the velocity magnitude plot (Fig. 9(a)), flow separation on the suction side occurs at a streamwise location of approximately $0.55c$. Estimating the location of the flow separation point is essential for determining the appropriate placement of the actuator on the suction side. According to the study by Ebrahimi and Hajipour [36], placing the actuator upstream of the separation point is more effective in mitigating flow separation.
Figure 9: Contours of baseline flow: (a) Velocity magnitude [m/s], (b) Turbulent kinetic energy [m2/s2], (c) Vorticity magnitude [1/s], and (d) Q-criterion [1/s2]. Source: Authors’ own work.
The TKE distribution (Fig. 9(b)) reveals the development of two free shear layers: one originating near the suction side separation point (approximately $0.55c$), and the other from the trailing edge. These shear layers evolve into a turbulent wake, as indicated by elevated energy levels. This dual-shear-layer structure is also clearly observed in the vorticity magnitude and Q-criterion contours (Figs. 9(c) and 9(d), respectively). The presence of these two shear layers provides the physical basis for employing the dual-point excitation strategy, which targets simultaneous stimulation of both regions [36]. Accordingly, in this study, the pressure side actuator is placed at the airfoil’s trailing edge.
@@ 5.2 Adaptive MPC
The adaptive MPC was evaluated at two target lift coefficient values, $C_{l}=1.6$ and $1.62$, to assess its effectiveness in controlling flow separation. These target values (set-points) were selected based on prior open-loop flow control studies, as well as on the present analysis of the system’s performance characteristics. First, a target $C_{l}$ of 1.6 was selected as the desired output. The objective was to adjust the excitation frequency of the plasma actuators in real time to achieve and maintain this lift coefficient. The control system dynamically responds to the airfoil’s transient aerodynamic characteristics to ensure that the $C_{l}$ reaches and stabilizes around the desired value of 1.6.
The graph presented in Fig. 10 illustrates the performance of the adaptive MPC system in achieving the desired lift coefficient. The top panel displays the time series of $C_{l}$ values along with its moving average, while the bottom panel shows the corresponding excitation frequencies applied during the control process. The lift coefficient stabilizes approximately 2.8 seconds after the controller is activated. According to the bottom panel, the actuator frequency converges to an excitation frequency of around 110 Hz. In terms of the normalized frequency, defined as $F^{+}=fc/U_{\infty}$, this corresponds to $F^{+}\approx 3$.
Figure 10: Desired $C_{l}$ of 1.6 using adaptive MPC and its corresponding excitation frequencies. Source: Authors’ own work.
To push the boundaries of the control system, aiming to achieve higher lift coefficients, a second target $C_{l}$ of 1.62 was selected. However, reaching such a high $C_{l}$ presents significant challenges for the adaptive MPC to stabilize $C_{l}$ around 1.62. As depicted in Fig. 11, the controller could not stabilize the $C_{l}$ at the desired value of 1.62. The larger fluctuations in the $C_{l}$ time series, compared with the previous case, reveal an unstable response as the controller struggles to regulate the set-point.
Unlike RL–based methods, which aim to maximize $C_{l}$ through reward-driven optimization, the MPC framework is designed to track a predefined set-point. Consequently, MPC continuously attempts to reach and hold the specified $C_{l}$. For the second target, $C_{l}=1.62$, this value lies near the physical limit of the dual plasma actuator configuration, making precise control difficult. Around $t=4.5$ s, MPC increases the excitation frequency in an effort to attain the target, but this adjustment shifts the system away from the optimal frequency and results in a reduction in lift. Since the desired $C_{l}$ exceeds what the actuators can sustainably produce, the controller fails to maintain the target, and the lift coefficient continues to oscillate below the set-point.
Figure 11: Desired $C_{l}$ of 1.62 using Adaptive MPC and its corresponding excitation frequencies. Source: Authors’ own work.
@@ 5.3 TDRL
Based on Fig. 12, the control system explored various frequencies before settling on 200 Hz as the optimal frequency for maximizing $C_{l}$. The top panel shows how $C_{l}$ evolves over time, while the bottom panel depicts the selected excitation frequencies throughout the process. By the end of the simulation, the system determined that 200 Hz was the most effective frequency for achieving the maximum $C_{l}$. Notably, after four decision-makings, the controller found the optimal frequency of 200 Hz, resulting in a mean $C_{l}$ of 1.619.
Figure 12: Exploring the optimal excitation frequency that maximizes $C_{l}$ using TDRL. Source: Authors’ own work.
@@ 5.4 DQL
As shown in Fig. 13, the DQL algorithm explored various frequencies before selecting 100 Hz as the optimal excitation frequency. Despite the fluctuations during the learning phase, the system eventually converged to 100 Hz as the most effective frequency for achieving a high and stable $C_{l}$. After 11 desicion makings, the controller found the optimal frequency of 100 Hz, resulting in a mean $C_{l}$ of 1.619.
Figure 13: Selection of the optimal excitation frequency for maximizing $C_{l}$ using DQL. Source: Authors’ own work.
@@ 5.5 Signal processing integrated with DQL
In this approach, DQL was enhanced by integrating signal processing techniques to improve the accuracy and robustness of the excitation frequency selection process. The objective was to refine the control system’s ability to identify the optimal frequency for maximizing the lift coefficient. According to Fig. 14, the system eventually selected 200 Hz as the optimal frequency. After 21 desicion makings, the controller found the optimal frequency of 200 Hz, resulting in a mean $C_{l}$ of 1.619.
Figure 14: Selection of the optimal excitation frequency for maximizing $C_{l}$ using signal processing integrated with DQL. Source: Authors’ own work.
The average $C_{l}$ for 100 Hz and 200 Hz are nearly identical, which made the decision-making process challenging for the network. The reason they are similar can be seen in the Fig. 15. For 100 Hz, there is one period of “on” and “off” states, while for 200 Hz, there are two periods. However, the first “on” phase for 200 Hz has a lower maximum than the “on” phase for 100 Hz, and the second “on” phase for 200 Hz reaches a higher maximum than 100 Hz. As a result, the average of two consecutive “on” states at 200 Hz closely matches the average of a single “on” state at 100 Hz.
This behavior suggests that the DQL produces similar average $C_{l}$ compared to the signal processing integrated with DQL and TDRL cases, making it harder for the network to differentiate between the two based purely on $C_{l}$ values. This similarity in average $C_{l}$ values helps explain why TDRL and signal processing integrated with DQL methods converged to 200 Hz, as their initial exploration behavior within the discrete action space, guided by early random sampling and the reward landscape, may have favored 200 Hz slightly, whereas DQL’s function approximator ultimately selected 100 Hz.
Figure 15: Comparison of results from different reinforcement learning methods for determining $C_{l}$. Source: Authors’ own work.
@@ 5.6 Comparative Analysis
Figure 16: Stabilization time and converged frequencies of different control methods. Source: Authors’ own work.
5.7 Aerodynamic Characteristics of the Controlled Flow
In this subsection, we examine the flow behavior under active control with dual-point excitation of plasma actuators. Fig. 17 illustrates the effect of plasma actuation on boundary layer velocity profiles within the plasma zones on both the suction and pressure sides of the airfoil. The horizontal axis represents the velocity magnitude normalized by the freestream velocity, while the vertical axis denotes the wall-normal position ($y_{n}$) normalized by the plasma zone height ($a$). Each plot includes twelve velocity profiles sampled at six streamwise locations within the plasma zones, depicted using a color gradient where darker shades correspond to greater $x/c$ positions. Red-shaded profiles indicate the plasma-on case with the DQL-integrated signal processing method, while blue-shaded profiles represent the plasma-off condition at the same streamwise locations.
Figure 17: Velocity profiles within the plasma actuator zones on the suction and pressure sides, shown with and without plasma actuation. Red and blue shading represent plasma-on and plasma-off conditions, respectively; darker shades indicate downstream streamwise locations. Source: Authors’ own work.
In the absence of plasma actuation, the boundary layer develops in a typical manner across the actuator regions. Notably, the velocity profiles on the suction side (Fig. 17(a)) exhibit lower fullness compared to those on the pressure side (Fig. 17(b)), due to the influence of an adverse pressure gradient on the suction side versus a favorable pressure gradient on the pressure side. Upon activation of the plasma actuators, an increase in profile fullness is observed at the upstream positions within the plasma zones, indicating momentum enhancement due to the added body force. Further downstream, the development of an ionic wind alters the boundary layer into a wall-jet-like profile with significantly higher momentum relative to the inactive case. Since the actuators operate in a periodic excitation mode, this results in the periodic formation and decay of the ionic wind. This mechanism plays a central role in generating instabilities in the flow field, which is a key feature of the active control strategy.
@@ 5.7 Aerodynamic Characteristics of the Controlled Flow
To compare the controlled flow field with the baseline flow (as shown in Fig. 9), Fig. 18 presents the contours of velocity magnitude, turbulent kinetic energy, vorticity magnitude, Q-criterion. These results are specifically provided for the signal processing integrated with DQL method. Since the flow fields for other control methods are nearly similar, their contours are not presented. This figure demonstrates that the proposed algorithms have successfully modified the flow field, including shear layers, wake, and separation points, compared to the baseline case.
@@
Figure 18: Contours of controlled flow using signal processing integrated with DQL method: (a) Velocity magnitude [m/s], (b) Turbulent kinetic energy [m2/s2], (c) Vorticity magnitude [1/s], and (d) Q-criterion [1/s2]. Source: Authors’ own work.
The velocity magnitude contours in Figs. 9(a) and 18(a) reveal a reduction in the extent of the separation zone under the influence of the flow control method. This observation is further supported by the reduced TKE levels in the suction-side separated shear layer of the controlled case compared to the baseline flow (Figs. 9(b) and 18(b)). Downstream of the trailing edge, the shear layer turns into discrete turbulent vortices. Additionally, the vorticity magnitude contours in Fig. 9(c) and 18(c) indicate that the separated shear layer on the airfoil’s suction side is inclined closer to the airfoil surface compared to the baseline case. This behavior promotes greater entrainment of high-energy free-stream fluid into the separated region on the suction side, thereby contributing to a reduction in the size of the separation zone. Furthermore, the coherent vorticity structures observed in the wake of the controlled flow suggest the presence of a periodic vortex-shedding pattern, attributed to the excitation of the shear layers.
The Q-criterion contours in Fig. 18(d) provide deeper insight into the influence of plasma actuator excitation on the development of the separated shear layer, compared to the baseline configuration shown in Fig. 9(d). Just downstream of the DBD plasma actuator (about the suction-side separation point) the controlled case exhibits alternating coherent regions of positive and negative Q values, indicating zones of elevated vorticity and strain rate, respectively. This pattern signifies the dominance of instabilities within the shear layer under plasma actuation. Moreover, the observed wavelength of these instabilities is synchronized with the excitation frequency of the actuator, a phenomenon known as lock-on, wherein shear layer dynamics resonate with external periodic forcing [35]. These synchronized instabilities accelerate the transition of the shear layer into turbulent, discrete vortices, thereby enhancing mixing between the separated region and the surrounding flow.
Fig. 19 compares the pressure coefficient ($C_{p}$) distribution along the airfoil surface for the baseline and controlled case. The locations of the two plasma actuators are indicated in red and blue, corresponding to the pressure and suction sides of the airfoil, respectively. On the pressure side, the control case exhibits slightly higher pressure values compared to the baseline. However, the overall difference in this region is not significant, as the flow remains largely attached. In contrast, the suction side shows a marked difference due to the applied flow control. From $x/c=0$ to approximately $x/c=0.65$, the controlled case exhibits a lower $C_{p}$ (i.e., higher suction) relative to the baseline. This increase in suction on the upper surface, coupled with the marginally elevated pressure on the lower surface, leads to a greater pressure differential across the airfoil and consequently results in an increased lift coefficient.
Figure 19: Comparison of $C_{p}$ distribution along $x/c$ between baseline flow and signal processing integrated with DQL. Source: Authors’ own work.